OpenStudy (jacobbirddog):

Fan And Medal! 5 questions! What is the solution set for the open sentence with the given replacement set? 13 – x = 7, {4, 5, 6, 7} A. 4 B. 5 C. 6 D. 7

OpenStudy (jacobbirddog):

What is the solution set for the open sentence with the given replacement set? 2x + 7 = 19, {4, 5, 6, 7} A. 4 B. 5 C. 6 D. 7

OpenStudy (jacobbirddog):

What is the solution set for the open sentence with the given replacement set? y2 – 3 = y + 9, {0, 1, 3, 4} A. 0 B. 1 C. 3 D. 4

OpenStudy (jacobbirddog):

Determine the width of the garden. Milo must put 50 feet of fencing around his rectangular garden. The length of the garden is one foot longer than twice the width. What is the solution set for 2w + 2(2w + 1) = 50 given the replacement set {6, 7, 8, 9}? A. 6 B. 7 C. 8 D. 9

OpenStudy (jacobbirddog):

Determine the rate at which Zayna traveled during the second part of her trip. Zayna drives for 3 hours at a rate of 44 miles per hour. Later, she drives for another 2 hours so that her toal distance traveled for the day is 240 miles. What is the solution set for 3(44) + 2r = 240 given the replacement set {52, 53, 54, 55}? A. 52 B. 53 C. 54 D. 55

OpenStudy (anonymous):

13-x=7 13-7=x x=6

OpenStudy (anonymous):

2x+7=19 2x=19-7 2x=12 x=6